From mar@roe.ac.uk Mon Feb 23 14:20:44 2015 Date: Sun, 22 Feb 2015 15:25:43 +0000 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , dmr@roe.ac.uk, Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , matt.jarvis@astro.ox.ac.uk, Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 13 February 2015 WFAU VDFS Science Archives progress meeting minutes, 13 February 2015 Present: NJC, RGM, ETWS, RPB, MAR, MSH, ACD, RSC Apologies: NCH, STV, KTN, AL DoNM: 10am Friday 27 February 2015 in the Villas meeting room, ACD in the chair (room is booked). Actions discharged this time: ----------------------------- ACTION: All to update the staff absences page. Done. ACTION ETWS/NJC: develop an automated tool to monitor progress in ingesting and curating surveys. Done and would be linked to the monitor web pages. ACTION: MAR check with Tom Shanks exactly what data are needed for the next ATLAS release. MAR had asked Tom about what his preferences were but no reply to date. Actions carried forward from 30/1/2015 meeting: ----------------------------------------------- ACTION: RPB (was MAR) to progress [was: 'have a look at'] the status of non-survey releases for registered projects. Continues. MAR had compiled a list of non-survey projects that could be released and sent it to RPB. RPB should progress these releases to completion. ACTION: RGM to investigate the size and content (how does it differ from DR9?) of SDSS DR10 to see if it is worth us hosting and whether it is necessary to keep all of the earlier releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. The part of this action that continues is that RGM had not yet established which early releases could safely be scrapped. RGM also noted that that DR11 and DR12 were now available and now need to be looked at. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this action still has low priority. ACTION: ACD to schedule a meeting to review progress with the various VVV brainstorming tasks from the previous meetings, RSC to chair. Continues. ACTION: ETWS/NCH purchase a new license for mxODBC. Continues. ETWS had raised this with NCH. ACTION: ACD to check with JAC whether they have dispatched all the tapes or there are more to be sent. Continues. ACTION: NJC noted that VISTA mosaicing crashes with normal filtering, which he should fix. Continues. ACTION: RPB to compile a Web page listing all the completed releases. Continues. Specific points and new actions: -------------------------------- Project management: Nothing to report. WFCAM, VISTA and VST updates: Nothing to report. Networking: There had been a short outage the previous week but other than that working well. WSA/VSA/OSA/GES Operations: ETWS reported that P92 VVV CU4 ingest were proceeding. The VVV DR3 tile-pawprinting had completed and CU2, CU3 and CU4 were now being run on the deeps. RPB noted that the VMC P93 was progressing following a CU6 fix. ATLAS (P91/DR2) had been restarted. Hardware and Systems: MSH reported that tenders were in the process of being sought for new hardware (DB node, webserver, processing node, NAS box). ACTION: NJC to produce roadmap for future VSA releases so as to plan for hardware requirements. Software: NJC reported working on further tweaks to the ESO release code and colour correction terms for the extinction maps. ETWS had been working on enhancements to the monitoring pages to show data release progress. Review of Deadlines: Progress towards forthcoming deadlines was reviewed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Survey Data Release: GESiDR2iDR3, which includes nightly spectra metadata and measurements. Updates to the VIDEO ESO releases. Non-survey Data Release: Nothing to report Astrogrid deployment, VO & Data Analysis services: MSH noted that the publishing registry has been periodically crashing but there were not many clues as to the cause. Miscellaneous: Following a request from AL, a twitter account had been set up to serve random cut-outs from the GPS/VVV Milky Way mosaic (@MilkyWayBot).